Circular No. 8475

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005W IN NGC 691 S. Nakano, Sumoto, Japan, reports the discovery by Yoji Hirose (Chi-

gasaki, Kanagawa-ken, Japan) of an apparent supernova (mag 15.2) on eight unfiltered CCD frames taken on Feb. 1.442 UT with a 0.35-m f/6.8 Schmidt-Cassegrain reflector. Nakano has measured the position of SN 2005W from one of Hirose’s images: α = 1^{h}50^{m}45^{s}.75, δ = +21^{o}45′35′′.6 (equinox 2000.0), which is 56′′ east and 1′′ south of the center of NGC 691. Hirose adds that nothing is visible at this location on frames taken by himself on Jan. 11, 21, and 24 (limiting mag 16.5–17) or on Digitized Sky Survey images. There is also a foreground star of mag 17.6 at position end figures 45^{s}.46, 38′′.6. W. Li, University of California, reports that a KAIT image taken on Feb. 2.18 shows the new object at mag 14.7, but nothing was visible at this location on a KAIT image from Jan. 23.20 (limiting mag 19.0); Li provides position end figures 45^{s}.77, 35′′.4 for SN 2005W.

SUPERNOVA 2005U IN ARP 299 M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithsonian Center

for Astrophysics, report that a spectrogram (range 350–740 nm) of SN 2005U (cf. IAUC 8473), obtained by P. Berlind on Feb. 1.51 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a type-II supernova, probably within a few weeks past explosion. The spectrum consists of a blue continuum and Balmer-line P-Cyg features with heavy host-galaxy light contamination. Adopting the NED recession velocity of 3088 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is ∼ 10000 km/s.

COMETS C/2004 Y9–Y11 AND C/2005 B2 (SOHO) Further to IAUC 8471, following are the initial observations of addi-

tional comets found on SOHO website images (Kreutz sungrazers except for C/2004 Y10).

Comet 2004 UT α_{2000} δ_{2000} Inst. F MPEC
C/2004 Y9 Dec. 28.034 18^{h}35.^{m}2 −24^{o}35′ C2  XG 2005-B62
C/2004 Y10 28.508 18 37.5 −22 16 C2 TH 2005-B62
C/2004 Y11  28.866 18 39.8 −25 20 C3/2 HS 2005-B62
Comet  2005 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2005 B2 Jan. 25.471 20^{h}46.^{m}7 −20^{o}26′ C3 KC 2005-B62
2005 February 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green